OpenStudy (anonymous):

Which shows the fractions in increasing order? seven-ninths, one-sixth, two-ninths, five-sixths one-sixth, two-ninths, seven-ninths, five-sixths one-sixth, two-ninths, five-sixths, seven-ninths one-sixth, five-sixths, two-ninths, seven-ninths

OpenStudy (anonymous):

are you allowed to use a calculator?

OpenStudy (anonymous):

there are several ways this question can be approached, One is you can put each in the calculator, find out the decimal and then arrange it.. in this case, 7/9 = 0.78 1/6 = 0.17 2/9 = 0.22 5/6 = 0.83 can you arrange it from here? 1/6 > 2/9 > ....

OpenStudy (anonymous):

you can also multiply the denominators with 6 or 9.. for example, \[\frac{ 7 * 6 }{ 9*6} = \frac{ 42 }{ 54 }\and \frac{ 1 * 9 }{ 9*6} = \frac{ 9 }{ 54 }\]

OpenStudy (anonymous):

do those respectively, then you should get 42/54, 9/54, 12/54, 45/54... Then arrange in increasing order by looking at the numerators. Because we are dealing with fractions, the highest numerator yields the lowest decimal.